# Practical hybrid decoding scheme for parity-encoded spin systems

### v3

@@ VI Conclusions
This paper proposed a practical decoding algorithm to correct readout errors tailored for the SLHZ model. Given the close connection between the SLHZ model and classical LDPC codes, classical decoding techniques for LDPC codes are expected to help exploit the latent potential of the SLHZ model. We proposed a simple BF algorithm based on iterative majority voting, a decoding algorithm for LDPC codes, which is much simpler than the standard BP algorithm. We demonstrated that the BF algorithm provides strong protection against i.i.d. noise for spin readouts of the SLHZ system and is as efficient as the BP algorithm. To study the tolerance to errors caused by a broader range of noise, we conducted a classical simulation of spin readouts on the SLHZ system using MCMC sampler. We found that BF decoding can correct leakage errors caused by dynamical and thermal excitations during MCMC sampling. Efficient decoding is possible if we reduce energy penalties, allowing thermal excitations to populate correctable leakage states more frequently. This is quite reasonable, given that stochastic sampling followed by classical error correction can be regarded as a two-stage hybrid decoding algorithm. Our observation suggests that the SLHZ model exhibits error correction against a broader range of noise models than the i.i.d. noise model when combined with post-readout BF decoding. Note that the BF algorithm does not suffer from the energy-barrier issue inherent to stochastic sampling. Thus, the two-stage hybrid decoding algorithm can be regarded as a practical implementation of the soft-annealing proposed by Sourlas [8]. Controlling the Lagrange weights in the first-stage sampler is crucial for efficiently obtaining the correct state. Special attention should be paid to the fact that the optimal Lagrange weights depend heavily on whether post-readout decoding is performed.
@@
Since our demonstration used readouts sampled stochastically by a classical method, specifically MCMC sampling, we are not necessarily convinced that the BF decoding is valid for readouts obtained through an SLHZ-based QA device. Further research is needed to assess how effectively the QA device can collaborate with BF decoding under realistic conditions. Nevertheless, we believe that most of our insights stem from the intrinsic nature of SLHZ model and BF decoding, and are applicable regardless of the stochastic sampling mechanism employed in the first stage. For example, if measurement error is the primary source of readout error of the SLHZ system, our BF decoding algorithm offers a straightforward solution to mitigate it. In addition, it would be reasonable to expect that a two-stage hybrid computation combining QA and post-readout BF decoding may address issues neither method can solve independently. Our research also emphasizes the importance of selecting appropriate decoding algorithms to exploit the potential of SLHZ systems.
@@
In this study, we tested only a small number of instances $K_{14}$ to investigate the performance and characteristics of the BF decoding; however, this is insufficient to draw a general conclusion. Of particular importance for future study is investigating the size dependence of the SLHZ model on the performance. In our discussion, we suggested that post-readout BF decoding may partially compensate for the poor spin-update properties of the SLHZ system. See the Appendix A in detail. In particular, it will be interesting to see how this compensation works well for a problem of arbitrary size. Albash et al. also pointed out another significant limitation of the SLHZ system: the strength of the energy penalty must grow with the problem size [24]. Introducing post-readout BF decoding may alleviate this limitation by reducing the penalty strength. Both these issues are important for achieving scalability. More research is needed to determine the efficacy of post-readout decoding.
@@ Acknowledgements.
I would like to thank Dr. T. Kadowaki at the Global Research and Development Center for Business by Quantum-AI technology (G-QuAT) and Prof. H. Nishimori at the Institute of Science Tokyo for their valuable comments and discussions. I also thank Dr. Masayuki Shirane of NEC Corporation/National Institute of Advanced Industrial Science and Technology for his continuous support. This paper is partly based on results obtained from a project, JPNP16007, commissioned by the New Energy and Industrial Technology Development Organization (NEDO), Japan.
@@
In Ref.[24], the authors present minimum weight decoding (MWD), which aims to find an error pattern $\boldsymbol{e}$ with minimum Hamming distance from the hard-decided readouts $\boldsymbol{r}$ and with the same syndrome vector as $\boldsymbol{s}^{(4)}(\boldsymbol{r})$. They directly estimated the optimal error pattern $\boldsymbol{e}^{*}\in\{\pm 1\}^{N_{v}}$ from the global ground state of the following Hamiltonian:
$H^{\mathrm{(MWD)}}(\boldsymbol{x})=-\sum_{i=1}^{N_{v}}x_{i}-\lambda\sum_{i=1}^{N_{c}}s_{i}^{(4)}(\boldsymbol{r})s_{i}^{(4)}(\boldsymbol{x}),$ (62)
@@
where $\boldsymbol{x}=(x_{1},\ldots,x_{N_{v}})$ is an arbitrary spin state, and $\boldsymbol{s}^{(4)}(\boldsymbol{x})=(s_{1}^{(4)}(\boldsymbol{x}),\ldots,s_{N_{c}}^{(4)}(\boldsymbol{x}))$ is the associated weight-4 syndrome vector. Here, $\lambda$ is assumed to be sufficiently large such that $\boldsymbol{x}$ minimizes all constraint terms. The second term in Eq.(62) forces $\boldsymbol{x}$ to satisfy $\boldsymbol{s}^{(4)}(\boldsymbol{x})=\boldsymbol{s}^{(4)}(\boldsymbol{r})$, and the first term minimizes the number of spins with $x_{i}=-1$. Let $C$ denote the set of $\boldsymbol{x}$ satisfying $\boldsymbol{s}^{(4)}(\boldsymbol{x})=\boldsymbol{s}^{(4)}(\boldsymbol{r})$. Then, the optimal estimate $\boldsymbol{e}^{*}$ is given by
$\boldsymbol{e}^{*}=\underset{\boldsymbol{x}\in C}{\arg\min}H^{\mathrm{(MWD)}}(\boldsymbol{x}).$ (63)
@@
Thus, $\boldsymbol{e}^{*}$ is the error pattern with the same syndrome vector and the smallest number of errors (i.e., the smallest number of elements $e_{k}$ with a value 1-1). Consequently, the state $\boldsymbol{z}^{*}$ estimated by the MWD is given by $\boldsymbol{z}^{*}=\boldsymbol{r}\circ\boldsymbol{e^{*}}$. The authors also related the MWD to finding the ground state of the following spin-glass Hamiltonian:
$H_{\mathrm{spin}}^{\mathrm{(MWD)}}(\boldsymbol{s})=-\sum_{i<j}g_{ij}s_{i}s_{j},$ (64)
@@
where a set of spin variables $\boldsymbol{g}=(g_{12},\ldots,g_{K-1\,K})\in\{\pm 1\}^{\tbinom{K}{2}}$ can be regarded as a gauge transformation. Once the ground state of Eq.(64) is found, one can deduce the optimal estimate $\boldsymbol{e}^{*}$ and obtain the estimate $\boldsymbol{z}^{*}$. The problem here is that the MWD requires the solution that minimizes the Hamiltonian defined in Eq. (64). This is equivalent to solving an instance of maximum 2-satisfiability, which is as hard as the original COP to be solved. Therefore, the MWD does not necessarily resolve the performance bottleneck for the SLHZ system. We think that MWD is challenging to apply in realistic scenarios when considering implementation in the near-term QA devices. In contrast, our BF decoding algorithm is an approximation, in which the current estimate $\boldsymbol{r}$ is iteratively updated to approach a better estimate $\boldsymbol{r}^{\prime}$. If the estimate $\boldsymbol{r}$ converges to a certain state within finite iterations, it gives a correct estimate $\boldsymbol{z}^{*}$ with a finite probability. Our BF algorithm can be executed on a deterministic Turing machine in polynomial time, since it completes with a finite number of iterations of matrix multiplication and the associated sign evaluation of every entry of the resultant matrix. Therefore, we believe our BF decoding algorithm is practical and realistic for implementation in near-term QA devices.
@@
We recognize that the performance boost of the MCMC-BF hybrid decoding at a vanishing penalty strength provides evidence supporting the inherent error-correcting capability of the SLHZ model. Our study suggested that controlling the penalty strength is crucial to exploiting the potential of the SLHZ model. Nonetheless, we do not claim that this performance boost is sufficient for the SLHZ model to outperform the ME model. We rather consider it essential that the SLHZ system be competitive with the ME model. For supplementary information, we have confirmed in our preliminary experiments that ME model shows no performance boost after post-readout MVD.
@@ Appendix B Supplementary result on experimental demonstration
In section IV, we demonstrated that the post-readout BF decoding algorithm can eliminate not only i.i.d. errors but also leakage errors due to dynamic and thermal excitations during the MCMC sampling process. However, when we imagine that the BF decoding is applied to the readout of actual QA devices rather than MCMC simulation, several types of errors may occur simultaneously. In this appendix, we assume i.i.d. errors in the readouts of MCMC sampler which might be a realistic model of readout errors for imperfect annealing devices, and investigate their impact on the performance of the MCMC-BF hybrid decoding. We present a result showing that the MCMC-BF decoder is robust to such readout errors to some extent.
Figure 12: Success probability for MCMC-BF decoder as a function of bit error rate $\varepsilon$ for the readouts of the MCMC sampler with four sample sizes. The sample size is expressed in a unit $N_{s}=\tbinom{K}{2}$, where $K=14$.
We simulate noisy readout $\hat{\boldsymbol{r}}^{\prime}$ by multiplying the error matrix $\hat{\boldsymbol{e}}$ generated by a pseudo-random number generator and the readout matrix $\hat{\boldsymbol{r}}$, i.e., $\hat{\boldsymbol{r}}^{\prime}=\hat{\boldsymbol{r}}\circ\hat{\boldsymbol{e}}$ , where $\hat{\boldsymbol{r}}$ was obtained from the same spin glass problem ($K_{14}$) as the main text. Subsequently, we applied the BF decoding to $\hat{\boldsymbol{r}}^{\prime}$, and investigated how the performance of the MCMC-BF hybrid decoder changes. To perform a statistical evaluation, we ran 1000 independent experiments. Figure12 shows the success probabilities for obtaining correct state when incorporating i.i.d. errors into the readouts of the MCMC sampler with four sample sizes. In this figure, we assumed that each spin flips with the bit error rate $\varepsilon$ $(0\leq\varepsilon\leq 0.5)$, and that the Lagrange weights are the values where the MCMC-BF hybrid decoder achieves maximum efficiency (corresponding to arrow B in Fig.9). As expected, the probabilities reduce as $\varepsilon$ increases and reach almost zero at $\varepsilon=0.5$. It was confirmed that decoding succeeds even when $\varepsilon\neq 0$, and, in particular, that the probability remains nearly unchanged in the region where $\varepsilon\leq 0.1$. This result suggests that the post-readout BF decoder can accommodate realistic noise in imperfect annealing devices.

### v4

@@ VI CONCLUSIONS
This paper has proposed a practical decoding algorithm to correct readout errors tailored for the SLHZ model. Given the close connection between the SLHZ model and classical LDPC codes, classical decoding techniques for LDPC codes are expected to help exploit the latent potential of the SLHZ model. We have proposed a simple BF algorithm based on iterative majority voting, a decoding algorithm for LDPC codes, which is much simpler than the standard BP algorithm. We have demonstrated that the BF algorithm provides strong protection against IID noise for spin readouts of the SLHZ system and is as efficient as the BP algorithm. To study the tolerance to errors caused by a broader range of noise, we have conducted a classical simulation of spin readouts on the SLHZ system using the MCMC sampler. We have found that BF decoding can correct leakage errors caused by dynamical and thermal excitations during MCMC sampling. Efficient decoding is possible if we reduce energy penalties, allowing thermal excitations to populate correctable leakage states more frequently. This is quite reasonable, given that stochastic sampling followed by classical error correction can be regarded as a two-stage hybrid decoding algorithm. Our observation suggests that the SLHZ model exhibits error correction against a broader range of noise models than the IID-noise model when combined with postreadout BF decoding. Note that the BF algorithm does not suffer from the energy-barrier issue inherent to stochastic sampling. Thus, the two-stage hybrid decoding algorithm can be regarded as a practical implementation of the soft-annealing proposed by Sourlas [8]. Controlling the Lagrange weights in the first-stage sampler is crucial for efficiently obtaining the correct state. Special attention should be paid to the fact that the optimal Lagrange weights depend heavily on whether postreadout decoding is performed.
@@
Since our demonstration has used readouts sampled stochastically by a classical method, specifically MCMC sampling, we are not necessarily convinced that the BF decoding is valid for readouts obtained through an SLHZ-based QA device. Further research is needed to assess how effectively the QA device can collaborate with BF decoding under realistic conditions. Nevertheless, we believe that most of our insights stem from the intrinsic nature of the SLHZ model and BF decoding and are applicable regardless of the stochastic sampling mechanism employed in the first stage. For example, if measurement error is the primary source of readout error of the SLHZ system, our BF decoding algorithm offers a straightforward solution to mitigate it. In addition, it would be reasonable to expect that a two-stage hybrid computation combining QA and postreadout BF decoding may address issues that neither method can solve independently. Our research also emphasizes the importance of selecting appropriate decoding algorithms to exploit the potential of SLHZ systems.
@@
In this study, we have tested only a small number of instances $K_{14}$ to investigate the performance and characteristics of the BF decoding; however, this is insufficient to draw a general conclusion. Of particular importance for future study is to investigate the size dependence of the SLHZ model on the performance. In our discussion, we have suggested that post-readout BF decoding may partially compensate for the poor spin-update properties of the SLHZ system (for details, see the Appendix A). In particular, it will be interesting to see whether this compensation works well for a problem of arbitrary size. Albash et al. have also pointed out another significant limitation of the SLHZ system: the strength of the energy penalty must grow with the problem size [24]. Introducing postreadout BF decoding may alleviate this limitation by reducing the penalty strength. Both these issues are important for achieving scalability. More research is needed to determine the efficacy of postreadout decoding.
@@ Acknowledgements.
I would like to thank Dr. T. Kadowaki at the Global Research and Development Center for Business by Quantum-AI technology (G-QuAT) and Professor H. Nishimori at the Institute of Science Tokyo for their valuable comments and discussions. I also thank Dr. Masayuki Shirane of NEC Corporation and the National Institute of Advanced Industrial Science and Technology for his continuous support. This paper is partly based on results obtained from a project, Grant No. JPNP16007, commissioned by the New Energy and Industrial Technology Development Organization (NEDO), Japan.
@@
In Ref.[24], the authors have present minimum-weight decoding (MWD), which aims to find an error pattern e with minimum Hamming distance from the hard-decided readouts r and with the same syndrome vector as $\textbf{s}^{(4)}(\textbf{r})$. They have directly estimated the optimal error pattern $\textbf{e}^{*}\in\{\pm 1\}^{N_{v}}$ from the global ground state of the following Hamiltonian:
$H^{\mathrm{(MWD)}}(\textbf{x})=-\sum_{i=1}^{N_{v}}x_{i}-\lambda\sum_{i=1}^{N_{c}}s_{i}^{(4)}(\textbf{r})s_{i}^{(4)}(\textbf{x}),$ (62)
@@
where $\textbf{x}=(x_{1},\ldots,x_{N_{v}})$ is an arbitrary spin state and $\textbf{s}^{(4)}(\textbf{x})=(s_{1}^{(4)}(\textbf{x}),\ldots,s_{N_{c}}^{(4)}(\textbf{x}))$ is the associated weight-4 syndrome vector. Here, $\lambda$ is assumed to be sufficiently large such that x minimizes all constraint terms. The second term in Eq.(62) forces x to satisfy $\textbf{s}^{(4)}(\textbf{x})=\textbf{s}^{(4)}(\textbf{r})$ and the first term minimizes the number of spins with $x_{i}=-1$. Let $C$ denote the set of x satisfying $\textbf{s}^{(4)}(\textbf{x})=\textbf{s}^{(4)}(\textbf{r})$. Then, the optimal estimate $\textbf{e}^{*}$ is given by
$\textbf{e}^{*}=\underset{\textbf{x}\in C}{\arg\min}H^{\mathrm{(MWD)}}(\textbf{x}).$ (63)
@@
Thus, $\textbf{e}^{*}$ is the error pattern with the same syndrome vector and the smallest number of errors (i.e., the smallest number of elements $e_{k}$ with a value 1-1). Consequently, the state $\textbf{z}^{*}$ estimated by the MWD is given by $\textbf{z}^{*}=\textbf{r}\circ\textbf{e}^{*}$. The authors have also related the MWD to finding the ground state of the following spin-glass Hamiltonian:
$H_{\mathrm{spin}}^{\mathrm{(MWD)}}(\textbf{s})=-\sum_{i<j}g_{ij}s_{i}s_{j},$ (64)
@@
where a set of spin variables $\textbf{g}=(g_{12},\ldots,g_{K-1\,K})\in\{\pm 1\}^{\tbinom{K}{2}}$ can be regarded as a gauge transformation. Once the ground state of Eq.(64) is found, one can deduce the optimal estimate $\textbf{e}^{*}$ and obtain the estimate $\textbf{z}^{*}$. The problem here is that the MWD requires the solution that minimizes the Hamiltonian defined in Eq. (64). This is equivalent to solving an instance of maximum 2-satisfiability, which is as hard as the original COP to be solved. Therefore, the MWD does not necessarily resolve the performance bottleneck for the SLHZ system. We think that MWD is challenging to apply in realistic scenarios when considering implementation in the near-term QA devices. In contrast, our BF decoding algorithm is an approximation, in which the current estimate r is iteratively updated to approach a better estimate $\textbf{r}^{\prime}$. If the estimate r converges to a certain state within finite iterations, it gives a correct estimate $\textbf{z}^{*}$ with a finite probability. Our BF algorithm can be executed on a deterministic Turing machine in polynomial time, since it completes with a finite number of iterations of matrix multiplication and the associated sign evaluation of every entry of the resultant matrix. Therefore, we believe that our BF decoding algorithm is practical and realistic for implementation in near-term QA devices.
@@
We recognize that the performance boost of the MCMC-BF hybrid decoding at a vanishing penalty strength provides evidence supporting the inherent error-correcting capability of the SLHZ model. Our study has suggested that controlling the penalty strength is crucial to exploiting the potential of the SLHZ model. Nonetheless, we do not claim that this performance boost is sufficient for the SLHZ model to outperform the ME model. Rather, we consider it essential that the SLHZ system be competitive with the ME model. For supplementary information, we have confirmed in our preliminary experiments that the ME model shows no performance boost after postreadout MVD.
@@ Appendix B SUPPLEMENTARY RESULT ON EXPERIMENTAL DEMONSTRATION
In section IV, we have demonstrated that the postreadout BF decoding algorithm can eliminate not only IID errors but also leakage errors due to dynamic and thermal excitations during the MCMC sampling process. However, when we imagine that the BF decoding is applied to the readout of actual QA devices rather than MCMC simulation, several types of errors may occur simultaneously. In this appendix, we assume IID errors in the readouts of the MCMC sampler which might be a realistic model of readout errors for imperfect annealing devices, and investigate their impact on the performance of the MCMC-BF hybrid decoding. We present a result showing that the MCMC-BF decoder is robust to such readout errors to some extent.
Figure 12: The success probability for the MCMC-BF decoder as a function of the bit error rate $\varepsilon$ for the readouts of the MCMC sampler with four sample sizes. The sample size is expressed in a unit $N_{s}=\tbinom{K}{2}$, where $K=14$.
We simulate noisy readout $\hat{\textbf{r}}^{\prime}$ by multiplying the error matrix $\hat{\textbf{e}}$ generated by a pseudorandom number generator and the readout matrix $\hat{\textbf{r}}$, i.e., $\hat{\textbf{r}}^{\prime}=\hat{\textbf{r}}\circ\hat{\textbf{e}}$ , where $\hat{\textbf{r}}$ has been obtained from the same spin-glass problem ($K_{14}$) as in the main text. Subsequently, we have applied the BF decoding to $\hat{\textbf{r}}^{\prime}$, and investigated how the performance of the MCMC-BF hybrid decoder changes. To perform a statistical evaluation, we have run 1000 independent experiments. In Fig.12, we show the success probabilities for obtaining the correct state when incorporating IID errors into the readouts of the MCMC sampler with four sample sizes. In this figure, we have assumed that each spin flips with the bit error rate $\varepsilon$ $(0\leq\varepsilon\leq 0.5)$, and that the Lagrange weights are the values at which the MCMC-BF hybrid decoder achieves maximum efficiency (corresponding to arrow B in Fig.9). As expected, the probabilities reduce as $\varepsilon$ increases and reach almost zero at $\varepsilon=0.5$. It has been confirmed that decoding succeeds even when $\varepsilon\neq 0$, and, in particular, that the probability remains nearly unchanged in the region in which $\varepsilon\leq 0.1$. This result suggests that the postreadout BF decoder can accommodate realistic noise in imperfect annealing devices.